# ReflecTool: Towards Reflection-Aware Tool-Augmented Clinical Agents

### v2

@@ Abstract
Yusheng Liao,♠,♢, Shuyang Jiang♣,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory {liao20160907,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 2.2 Clinical Toolbox
ReflecTool (Iterative Refinement, n=2) 163.79 160.83 121.97 151.65 148.60 149.37
@@ 5 Analysis
5.1 Tool Usage Error
Although we categorized tools based on the capability dimensions of the model (as shown in Table 6), using tools from other dimensions is not necessarily incorrect. For example, when addressing questions in the Knowledge dimension, it is reasonable to employ the NER model from the Data Understanding dimension to extract specialized terms, which can then be used to query a knowledge graph more effectively. However, there are still some cases of incorrect tool usage for specific tools: (1) invoking an MLLM without a medical image input, (2) using SQLCoder and DBManual without a database input, and (3) employing LongDocRAG without document file input. Based on these points, we analyze the behavior of different methods in selecting tools. As shown in Table 4, ReflecTool can significantly reduce the tool selection error on both step-level and task-level. The results support that the tool-wise reflection mechanism can improve agents’ ability to use appropriate domain tools. Besides, we also analyze the parameter error of tool usage in Appendix C.1.
Tool Selection Error $\downarrow$
Methods Step-Level Task-Level
ReAct (Yao et al., 2023) 1.44 4.03
Reflexion (Gou et al., 2024) 0.92 1.67
CRITIC (Shinn et al., 2023) 0.24 0.33
ReflecTool (Iterative Refinement) 0.06 0.15
ReflecTool (Candidate Selection) 0.02 0.08
@@ B.1.4 Data Understanding
emrQA (Pampari et al., 2018) is a large-scale question-answering (QA) dataset specifically designed for clinical notes. It is constructed by leveraging existing expert annotations from the i2b2666https://www.i2b2.org/NLP/DataSets/ datasets, resulting in a dataset with over 1 million question-logical form pairs and more than 400,000 question-answer evidence pairs. emrQA aims to support the development of QA systems capable of understanding complex clinical narratives and providing answers based on longitudinal patient records. Here, we derive the emrQA dataset from Huggingface [^7] https://huggingface.co/datasets/Eladio/emrqa-msquad.
@@ B.2.1 Knowledge Tools
Following the implementation of Xiong et al. (2024), the knowledge base in our method is composed of PubMed [^9] https://pubmed.ncbi.nlm.nih.gov/, StatPearls [^10] https://www.statpearls.com/, and Textbooks (Jin et al., 2021). We adopt BM25111111https://github.com/facebookresearch/faiss (Robertson et al., 2009) as the retriever to search the information from the medical knowledge base.

### v3

@@ Abstract
Yusheng Liao♠,♢, Shuyang Jiang♣,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Shanghai Artificial Intelligence Laboratory Fudan University {liao20160907,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn shuyangjiang23@m.fudan.edu.cn
@@ 2.2 Clinical Toolbox
ReflecTool (Iterative Refinement, n=2) 163.79 160.83 121.97 151.65 48.60 149.37
@@ 5.4 Tool Distribution in Trajectory
Tool Selection Error $\downarrow$
Methods Step-Level Task-Level
ReAct (Yao et al., 2023) 1.44 4.03
Reflexion (Gou et al., 2024) 0.92 1.67
CRITIC (Shinn et al., 2023) 0.24 0.33
ReflecTool (Iterative Refinement) 0.06 0.15
ReflecTool (Candidate Selection) 0.02 0.08
Methods Tools Error Rate Task Error Rate
SQLCoder DBManual LongDocRAG Avg. MedQA MMLU BioASQ SLAKE MedCalc EHRSQL MedMen LongHealth EMRQA Avg.
ReAct 11.99 0.00 5.86 5.95 0.00 0.00 0.00 0.08 0.00 0.07 0.08 8.71 12.35 2.36
CRITIC 1.74 20.51 1.19 7.81 0.00 0.00 0.00 0.00 0.03 0.07 0.00 5.12 0.38 0.62
Reflexion 3.13 2.05 34.05 13.08 0.02 0.02 0.04 0.00 0.73 0.25 0.00 1.30 8.79 1.24
ReflecTool (Iterative Refinement) 0.02 0.00 6.40 2.14 0.00 0.00 0.00 0.00 0.21 0.02 0.00 0.00 0.43 0.07
ReflecTool (Candidate Selection) 0.07 0.00 1.95 0.67 0.00 0.02 0.04 0.00 0.11 0.02 0.00 0.00 0.00 0.02
Table 6: Tool selection error rates with specific tool and task. Each value indicates the percentage of incorrect tool selections made by the model when using that tool or performing that task. Columns with an entirely zero error rate have been omitted.
5.5 Tool Usage Error
Although we categorized tools based on the capability dimensions of the model (as shown in Table 9), using tools from other dimensions is not necessarily incorrect. For example, when addressing questions in the Knowledge dimension, it is reasonable to employ the NER model from the Data Understanding dimension to extract specialized terms, which can then be used to query a knowledge graph more effectively. However, there are still some cases of incorrect tool usage for specific tools: (1) invoking an MLLM without a medical image input, (2) using SQLCoder and DBManual without a database input, and (3) employing LongDocRAG without document file input. Based on these points, we analyze the behavior of different methods in selecting tools. As shown in Table 5, ReflecTool can significantly reduce the tool selection error on both step-level and task-level. The results support that the tool-wise reflection mechanism can improve agents’ ability to use appropriate domain tools.
To further understand the distribution of tool selection errors, we separately counted the instances of invocation errors for different methods across tools and datasets. The results are shown in Table 6. From the perspective of tool types, it can be seen that errors mainly occur with three types of tools: SQLCoder, DBManual, and LongDocRAG. These tools require specific inputs to function, namely databases and uploaded files. Therefore, it is likely that the model misuses these tools when the corresponding inputs are missing. From the perspective of task types, errors are predominantly concentrated in LongHealth, EMRQA, and EHR tasks. These tasks share common points of confusion because their input context is similar, but with different formats. For instance, both LongHealth and EMRQA involve contextual question answering, but LongHealth, due to its ultra-long context, must be uploaded as a file. This again confirms that the model’s ability to match other information and appropriate tool usage, aside from image inputs, is relatively weak. In such scenarios, ReflecTool can effectively identify the relationship between tools and their corresponding input information modalities, thereby preventing such errors. Besides, we also analyze the parameter error of tool usage in Appendix C.1.
Methods Avg.
ReAct (Yao et al., 2023) 55.85
Reflexion (Gou et al., 2024) 58.78
CRITIC (Gou et al., 2024) 57.29
ReflecTool (Iterative Refinement) 62.26
ReflecTool (Candidates Selection) 60.72
Table 7: Performance of agent-based methods with Llama3-70B-Instruction as the backbone.
Methods Qwen2-7b Qwen2-72b
ReAct (Yao et al., 2023) 11.01 20.04
CRITIC (Gou et al., 2024) 4.33 17.87
Reflexion (Gou et al., 2024) 12.42 57.27
ReflecTool (Iterative Refinement) 11.95 47.90
ReflecTool (Candidate Selection) 11.26 28.56
Table 8: Runtime (seconds per sample) of the agent-based method in single-sample tests. Note that lower values indicate higher agent efficiency.
5.6 Performance with Different Backbones
During our experiments, we found that the instruction-following capabilities of the Llama series models were inferior to those of the Qwen2 series models, often resulting in formatting errors. Therefore, we prioritized the Qwen series to implement our methodology. After carefully adjusting the prompts, we also tested the performance of Agent-based methods on ablation subsets using the Llama3-70B-Instruction [^3] https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct model. The results in Table 7 showed that the proposed ReflecTool method is also effective on the Llama3-70B-Instruction model, demonstrating the generalization ability of this method across different models and thus indicating better application potential.
5.7 Time Cost Analysis
A practical method should enhance performance without incurring excessive resource consumption. It can be observed from Table 8 that for the strong baseline Reflxion, ReflecTool consumes less time on both 7b and 70b models. It’s noteworthy that Candidates Selection is faster than Iteration Refinement, which is due to the higher degree of parallelism in the former. Besides, the time consumption of CRITIC is less for it does not utilize too many tools to solve tasks in many cases. Given that tool invocation requires more time compared to decision-making by large models (for instance, MedRAG requires retrieval, and UMLS needs internet access), this leads to lower time consumption but poor performance. In summary, ReflecTool proves to be highly efficient from the perspective of resource consumption.
@@ 6 Conclusions
Acknowledgements
This work was supported by the National Key R&D Program of China (No. 2022ZD0162101) and STCSM (No. 22DZ2229005)
@@ B.1.4 Data Understanding
emrQA (Pampari et al., 2018) is a large-scale question-answering (QA) dataset specifically designed for clinical notes. It is constructed by leveraging existing expert annotations from the i2b2777https://www.i2b2.org/NLP/DataSets/ datasets, resulting in a dataset with over 1 million question-logical form pairs and more than 400,000 question-answer evidence pairs. emrQA aims to support the development of QA systems capable of understanding complex clinical narratives and providing answers based on longitudinal patient records. Here, we derive the emrQA dataset from Huggingface [^8] https://huggingface.co/datasets/Eladio/emrqa-msquad.
@@ B.2.1 Knowledge Tools
Following the implementation of Xiong et al. (2024), the knowledge base in our method is composed of PubMed [^10] https://pubmed.ncbi.nlm.nih.gov/, StatPearls [^11] https://www.statpearls.com/, and Textbooks (Jin et al., 2021). We adopt BM25121212https://github.com/facebookresearch/faiss (Robertson et al., 2009) as the retriever to search the information from the medical knowledge base.